# Frame Interpolation with Consecutive Brownian Bridge Diffusion

### v3

@@ Abstract.
Figure 1. Qualitative Comparison of our proposed method and SOTAs. Our method generates clear interpolated images, while recent SOTAs generate blurred or overlaid results. Images within blue boxes are displayed for detailed comparisons, and red circles highlight our performance. Examples are chosen from SNU-FILM (Choi et al., 2020) extreme subset. More visualization results are shown in Appendix C.2
@@ 1. Introduction
Current works in VFI can be divided into two folds in terms of methodologies: flow-based methods (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021; Dutta et al., 2022; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023) and kernel-based methods (Chen et al., 2021; Cheng and Chen, 2020; Lee et al., 2020; Niklaus et al., 2017a, b; Shi et al., 2021; Dai et al., 2017). Flow-based methods take advantages of optical flow estimations to warp images and features to synthesize intermediate frames. Instead of relying on optical flows, kernel-based methods predict convolution kernels for pixels in the neighboring frames. Recent advances in flow estimation (Huang et al., 2022a; Weinzaepfel et al., 2023; Hui et al., 2018; Huang et al., 2022c; Ilg et al., 2017; Teed and Deng, 2020; Sun et al., 2018) make it more popular to adopt flow-based methods in VFI.
@@
    The formulation of diffusion models results in a large cumulative variance (the variance accumulated during sampling) of generated latent representations. The sampling process starts with standard Gaussian noise and adds small Gaussian noise to the denoised output at each step. Noises are added up to a large cumulative variance when images are generated. Though such a variance is beneficial to diversity (i.e. repeated sampling results in different outputs), VFI requires that repeated sampling returns identical results, which is the ground truth intermediate frame. Therefore, a small cumulative variance is preferred in VFI. The relation of the cumulative variance and diversity is supported by the fact that DDIM (Song et al., 2021a) tends to generate relatively deterministic images than DDPM (Ho et al., 2020) because DDIM removes small noises at each sampling step. In Section 3.4, we show that our method has a much lower cumulative variance.
@@
To tackle these challenges, we propose a consecutive Brownian Bridge Diffusion (in latent space) that transits among three deterministic endpoints for VFI. This method results in a much smaller cumulative variance, achieving a better estimation of the ground truth inputs. We also provide a novel method to analyze the LDM-based VFI methods: by analyzing the gap between quantitative metrics of the outputs from the autoencdoer and the outputs from the diffusion model + decoder (we name this as ground truth estimation), it is easier to figure out the specific directions of improvement. Moreover, we take advantage of flow estimation and refinement methods in recent literature (Lu et al., 2022a) to improve the autoencoder. The feature pyramids from neighboring frames are warped based on estimated optical flows, aiming to alleviate the issues of reconstructing overlaid images. In experiments, our method improves by a large margin when the autoencoder is improved and achieves state-of-the-art performance. Our contribution can be summarized in three parts:
  • We propose a novel consecutive Brownian Bridge diffusion model for VFI and justify its advantages over traditional diffusion models: lower cumulative variance and better ground truth estimation capability. Additionally, we provide a cleaner formulation of Brownian Bridges and also propose the loss weightng for Brownian Bridges.
@@
  • Through extensive experiments, we validate the effectiveness of our method. Our method estimates the ground truth better than traditional diffusion with conditional generation. Moreover, the performance of our method improves when the autoencoder improves and achieves state-of-the-art performance with a simple yet effective autoencoder, indicating its strong potential in VFI.
@@ 3.1. Preliminaries
(9) $q(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{y})=\mathcal{N}\left(\mathbf{x}_{t};(1-m_{t})\mathbf{x}_{0}+m_{t}\mathbf{y},\delta_{t}\right),\vspace{-2mm}$
@@
where $\text{where }\tilde{\mathcal{\mu}}_{t}=c_{xt}\mathbf{x}_{t}+c_{yt}y+c_{\epsilon t}(m_{t}(\mathbf{y}-\mathbf{x}_{0})+\sqrt{\delta_{t}}\epsilon),\vspace{-3mm}$
$c_{xt}=\frac{\delta_{t-1}}{\delta_{t}}\frac{1-m_{t}}{1-m_{t-1}}+\frac{\delta_{t|t-1}}{\delta_{t}}(1-m_{t}),\vspace{-1mm}$
$c_{yt}=m_{t-1}-m_{t}\frac{1-m_{t}}{1-m_{t-1}}\frac{\delta_{t-1}}{\delta_{t}},\vspace{-1mm}$
$c_{\epsilon t}=(1-m_{t-1})\frac{\delta_{t|t-1}}{\delta_{t}},\vspace{-1mm}$
@@ 3.3. Autoencoder
Decoding with Warped Features. LDMVFI (Danier et al., 2024) apply cross-attention (Vaswani et al., 2017) to up-sampled $\mathbf{\hat{x}}$ and $f_{x}^{k},f_{y}^{k}$. However, this does not explicitly deal with motion changes, and therefore LDMVFI usually produces overlaid results as shown in Fig. 1. Therefore, we estimate optical flows from $I_{n}$ to $I_{0},I_{1}$ and apply backward warping to the feature pyramids to tackle this problem. Suppose $\hat{x}$ is generated by our consecutive Brownian Bridge diffusion, and it is up-sampled to $h^{k}$ where $k$ denotes the down-sampling factor compared to the original image. Then, we apply $CA\left(h^{k},Cat(w(f_{y}^{k}),w(f_{z}^{k}))\right)$ for $k>1$ to fuse the latent representation $h^{k}$ and feature pyramids $f_{y}^{k}$ and $f_{z}^{k}$, where $CA(\cdot,\cdot)$, $Cat(\cdot,\cdot)$, and $w(\cdot)$ denotes cross attention, channel-wise concatenation, and backward warping with estimated optical flows respectively. Finally, we apply convolution layers to $h^{1}$ to predict soft mask $H$ and residual $\delta$. The interpolation output is $\hat{I}_{n}=H*w(I_{0})+(1-H)*w(I_{1})+\delta$, where * holds for Hadamard product, and $\hat{I}_{n}$ is the reconstructed image. The detailed illustration of the architecture is shown in Figure 3. The VQ layer is connected with the encoder during training, but it is disconnected from the encoder and receives the sampled latent representation from the diffusion model.
@@ 4.1. Implementations
Table 2. Ablation studies of autoencoder and ground truth estimation. + GT means we input ground truth $\mathbf{x}$ to the decoder. + BB indicates our consecutive Brownian Bridge diffusion trained with autoencoder of LDMVFI. With our consecutive Brownian Bridge, the interpolated frame has highly similar results with the decoded frame from ground truth latent representation, indicating the strong ground truth estimation capability. Our autoencoder also has better performance than LDMVFI (Danier et al., 2024).
@@ 4.2. Datasets and Evaluation Metrics
Evaluation Metrics. Recent works (Danier et al., 2024, 2022; Zhang et al., 2018) reveal that PSNR and SSIM (Wang et al., 2004) are sometimes unreliable because they have relatively low correlation with humans’ visual judgments. However, learning-based metrics such as FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and FloLPIPS (Danier et al., 2022) are shown to have a higher correlation with humans’ visual judgments in (Danier et al., 2024; Zhang et al., 2018). Moreover, we also experimentally find such inconsistencies between PSNR/SSIM and visual quality, which will be discussed in Section 4.3. Therefore, we select FID, LPIPS, and FloLPIPS as our main evaluation metrics. LPIPS and FID measure distances in the space of deep learning features. FloLPIPS is based on LPIPS but takes the motion in the frames into consideration. The results in PSNR/SSIM are included in Appendix C.1
@@ 4.3. Experimental Results
Figure 4. The reconstruction quality of our autoencoder and LDMVFI’s. Images are cropped with green boxes for detailed comparisons. Red circles highlight our performance, where LDMVFI usually outputs overlaid images.
@@
Qualitative Results. We have provided visual comparisons of our method and recent SOTAs in Figure 1. Our method achieves better visual quality because we have clearer details such as dog skins, cloth with folds, and fences with nets. However, UPR-Net (Jin et al., 2023) achieves better PSNR/SSIM in all the cropped regions ($5-10\%$ better) than ours, which is highly inconsistent with the visual quality. Moreover, as mentioned in Section 1, we find that the autoencoder in (Danier et al., 2024) usually reconstructs overlaid images, and therefore we propose a new method of reconstruction. We provide examples to visualize the reconstruction results with our autoencoder and LDMVFI’s for comparison, shown in Figure 4. All examples are from SNU-FILM extreme (Choi et al., 2020), which contains relatively large motion changes in neighboring frames.
@@ 4.4. Ablation Studies
As we discussed in Section 3.2, latent-diffusion-based VFI is broken down into two stages, so we have a novel method to analyze the entire model. We conduct an ablation study on the ground truth estimation capability of our consecutive Brownian Bridge diffusion. We compare the evaluation results of decoded images with diffusion-generated latent representation $\hat{\mathbf{x}}$ and ground truth $\mathbf{x}$, which is encoded $I_{n}$. The results are shown in Table 2. It is important to note that, fixing inputs as the ground truth, our autoencoder achieves a stronger performance than the autoencoder in LDMVFI (Danier et al., 2024), indicating the effectiveness of our autoencoder. Also, fixing the autoencoder, our consecutive Brownian Bridge diffusion achieves almost identical performance with the ground truth, indicating its strong capability of ground truth estimation. However, the conditional generation model in LDMVFI (Danier et al., 2024) usually underperforms the autoencoder with ground truth inputs. Therefore, our method has a stronger ability in both the autoencoder and ground truth estimation stages. More ablation studies are provided in Appendix C.3
@@ 5. Conclusion
In this study, we propose our consecutive Brownian Bridge diffusion Model that better estimates the ground truth latent representation due to its low cumulative variance. We justify its effectiveness with extensive experiments in a wide range of datasets. Our method improves when the autoencoder is improved and achieves state-of-the-art performance with a simple yet effective design of the autoencoder, demonstrating its strong potential in the VFI task as a carefully designed autoencoder could potentially boost the performance by a large margin. In addition, we propose a novel method to analyze LDM-based VFI, providing insights for future research. Therefore, we believe our work will provide unique research directions and insights for diffusion-based video frame interpolation.
@@ C.3. Ablation Studies
Consecutive Brownian Bridge Diffusion. We provide some visualizations to support the result of our Consecutive Brownian Bridge Diffusion generates higher quality results with the same autoencoder than LDMVFI (Danier et al., 2024), shown in Figure 7.

### v4

@@ Abstract.
Figure 1. Qualitative Comparison of our proposed method and SOTAs. Our method generates clear interpolated images, while recent SOTAs generate blurred or overlaid results. In our method, the generated images have clearer dog skins (first row), clearer cloth with folds (second row), and clearer fences with nets and high-quality shoes (third row). Images within blue boxes are displayed for detailed comparisons, and red circles highlight our performance. Examples are chosen from SNU-FILM (Choi et al., 2020) extreme subset. More visualization results are shown in Appendix C.2
@@ 1. Introduction
Current works in VFI can be divided into two folds in terms of methodologies: flow-based methods (Plack et al., 2023; Lu et al., 2022a; Jin et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Siyao et al., 2021; Hu et al., 2022; Niklaus and Liu, 2018; Choi et al., 2021; Dutta et al., 2022; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023) and kernel-based methods (Chen et al., 2021; Cheng and Chen, 2020; Lee et al., 2020; Niklaus et al., 2017a, b; Shi et al., 2021; Dai et al., 2017). Flow-based methods compute flows in the neighboring frames and forward warp neighboring images and features (Hu et al., 2022; Niklaus and Liu, 2018, 2020; Jin et al., 2023; Siyao et al., 2021) or estimate flows from the intermediate frame to neighboring frames and backward warp neighboring frames and features (Lu et al., 2022a; Plack et al., 2023; Argaw and Kweon, 2022; Huang et al., 2022b; Dutta et al., 2022; Choi et al., 2021; Li et al., 2023b; Park et al., 2023; Zhang et al., 2023). Instead of relying on optical flows, kernel-based methods predict convolution kernels for pixels in the neighboring frames. Recent advances in flow estimation (Huang et al., 2022a; Weinzaepfel et al., 2023; Hui et al., 2018; Huang et al., 2022c; Ilg et al., 2017; Teed and Deng, 2020; Sun et al., 2018) make it more popular to adopt flow-based methods in VFI.
@@
    The formulation of diffusion models results in a large cumulative variance (the variance accumulated during sampling) of generated latent representations. The sampling process starts with standard Gaussian noise and adds small Gaussian noise to the denoised output at each step. Noises are added up to a large cumulative variance when images are generated. Though such a variance is beneficial to diversity (i.e. repeated sampling results in different outputs), VFI requires that repeated sampling returns identical results, which is the ground truth intermediate frame. Therefore, a small cumulative variance is preferred in VFI. The relation of the cumulative variance and diversity is supported by the fact that DDIM (Song et al., 2021a) tends to generate relatively deterministic images than DDPM (Ho et al., 2020) because DDIM removes small noises at each sampling step. LDMVFI (Danier et al., 2024) uses conditional generation as guidance, but this does not change the nature of large cumulative variance. In Section 3.4, we show that our method has a much lower cumulative variance.
@@
To tackle these challenges, we propose a consecutive Brownian Bridge Diffusion (in latent space) that transits among three deterministic endpoints for VFI. This method results in a much smaller cumulative variance, achieving a better estimation of the ground truth inputs. We also provide a novel method to analyze the LDM-based VFI methods: by analyzing the gap between quantitative metrics of the outputs from the autoencoder and the outputs from the diffusion model + decoder (we name this as ground truth estimation), it is easier to figure out the specific directions of improvement: whether to improve the autoencoder or the diffusion model. Moreover, we take advantage of flow estimation and refinement methods in recent literature (Lu et al., 2022a) to improve the autoencoder. The feature pyramids from neighboring frames are warped based on estimated optical flows, aiming to alleviate the issues of reconstructing overlaid images. In experiments, our method improves by a large margin when the autoencoder is improved and achieves state-of-the-art performance. Our contribution can be summarized in three parts:
  • We propose a novel consecutive Brownian Bridge diffusion model for VFI and justify its advantages over traditional diffusion models: lower cumulative variance and better ground truth estimation capability. Additionally, we provide a cleaner formulation of Brownian Bridges and also propose the loss weighting for our Consecutive Brownian Bridges.
@@
  • Through extensive experiments, we validate the effectiveness of our method. Our method estimates the ground truth better than traditional diffusion with conditional generation (LDMVFI (Danier et al., 2024)). Moreover, the performance of our method improves when the autoencoder improves and achieves state-of-the-art performance with a simple yet effective autoencoder, indicating its strong potential in VFI.
@@ 3.1. Preliminaries
(9) $q(\mathbf{x}_{t}|\mathbf{x}_{0},\mathbf{y})=\mathcal{N}\left(\mathbf{x}_{t};(1-m_{t})\mathbf{x}_{0}+m_{t}\mathbf{y},\delta_{t}\right),$
@@
where $\text{where }\tilde{\mathcal{\mu}}_{t}=c_{xt}\mathbf{x}_{t}+c_{yt}y+c_{\epsilon t}(m_{t}(\mathbf{y}-\mathbf{x}_{0})+\sqrt{\delta_{t}}\epsilon),$
$c_{xt}=\frac{\delta_{t-1}}{\delta_{t}}\frac{1-m_{t}}{1-m_{t-1}}+\frac{\delta_{t|t-1}}{\delta_{t}}(1-m_{t}),$
$c_{yt}=m_{t-1}-m_{t}\frac{1-m_{t}}{1-m_{t-1}}\frac{\delta_{t-1}}{\delta_{t}},$
$c_{\epsilon t}=(1-m_{t-1})\frac{\delta_{t|t-1}}{\delta_{t}},$
@@ 3.3. Autoencoder
Decoding with Warped Features. LDMVFI (Danier et al., 2024) apply cross-attention (Vaswani et al., 2017) to up-sampled $\mathbf{\hat{x}}$ and $f_{x}^{k},f_{y}^{k}$. However, this does not explicitly deal with motion changes, and therefore LDMVFI usually produces overlaid results as shown in Figure 1. Therefore, we estimate optical flows from $I_{n}$ to $I_{0},I_{1}$ and apply backward warping to the feature pyramids to tackle this problem. Suppose $\hat{x}$ is generated by our consecutive Brownian Bridge diffusion, and it is up-sampled to $h^{k}$ where $k$ denotes the down-sampling factor compared to the original image. Then, we apply $CA\left(h^{k},Cat(w(f_{y}^{k}),w(f_{z}^{k}))\right)$ for $k>1$ to fuse the latent representation $h^{k}$ and feature pyramids $f_{y}^{k}$ and $f_{z}^{k}$, where $CA(\cdot,\cdot)$, $Cat(\cdot,\cdot)$, and $w(\cdot)$ denotes cross attention, channel-wise concatenation, and backward warping with estimated optical flows respectively. Finally, we apply convolution layers to $h^{1}$ to predict soft mask $H$ and residual $\delta$. The interpolation output is $\hat{I}_{n}=H*w(I_{0})+(1-H)*w(I_{1})+\delta$, where * holds for Hadamard product, and $\hat{I}_{n}$ is the reconstructed image. The detailed illustration of the architecture is shown in Figure 3. The VQ layer is connected with the encoder during training, but it is disconnected from the encoder and receives the sampled latent representation from the diffusion model.
@@ 4.1. Implementations
Table 2. Ablation studies of autoencoder and ground truth estimation. + GT means we input ground truth x to the decoder part of autoencoder. + BB indicates our consecutive Brownian Bridge diffusion trained with autoencoder of LDMVFI. With our consecutive Brownian Bridge diffusion, the interpolated frame has almost the same performance as the interpolated frame with ground truth latent representation, indicating the strong ground truth estimation capability Our autoencoder also has better performance than LDMVFI (Danier et al., 2024).
@@ 4.2. Datasets and Evaluation Metrics
Evaluation Metrics. Recent works (Danier et al., 2024, 2022; Zhang et al., 2018) reveal that PSNR and SSIM (Wang et al., 2004) are sometimes unreliable because they have relatively lower correlation with humans’ visual judgments. However, learning-based metrics such as FID (Heusel et al., 2017), LPIPS (Zhang et al., 2018), and FloLPIPS (Danier et al., 2022) are shown to have a higher correlation with humans’ visual judgments in (Danier et al., 2024; Zhang et al., 2018). Moreover, we also experimentally find such inconsistencies between PSNR/SSIM and visual quality, which will be discussed in Section 4.3. Therefore, we select FID, LPIPS, and FloLPIPS as our main evaluation metrics. LPIPS and FID measure similarities or distances in the latent space of deep learning models. FloLPIPS is based on LPIPS but takes the motion change among three frames into consideration. The results in PSNR/SSIM are included in Appendix C.1.
@@ 4.3. Experimental Results
Qualitative Results. In Table 1, our consecutive Brownian Bridge diffusion with the autoencoder in LDMVFI (Danier et al., 2024) (denoted as $\dagger$) generally achieves better quantitative results than LDMVFI, showing our method is effective. We include qualitative visualization in Figure 5 to support this result. Moreover, as mentioned in Section 1, we find that the autoencoder in (Danier et al., 2024) usually reconstructs overlaid images, and therefore we propose a new method of reconstruction. We provide examples to visualize the reconstruction results with our autoencoder and LDMVFI’s autoencoder for comparison, shown in Figure 4. All examples are from SNU-FILM extreme (Choi et al., 2020), which contains relatively large motion changes in neighboring frames.
Figure 4. The reconstruction quality of our autoencoder and LDMVFI’s autoencoder (decoding with ground truth latent representation x). Images are cropped with green boxes for detailed comparisons. Red circles highlight the details where our method achieves better performance. LDMVFI usually outputs overlaid images while our method does not.
We have provided some visual comparisons of our method and recent SOTAs in Figure 1. Our method achieves better visual quality because we have clearer details such as dog skins, cloth with folds, and fences with nets. However, UPR-Net (Jin et al., 2023) achieves better PSNR/SSIM in all the cropped regions ($5-10\%$ better) than ours, which is highly inconsistent with the visual quality.
@@ 4.4. Ablation Studies
As we discussed in Section 3.2, latent-diffusion-based VFI is broken down into two stages, so we have a novel method to analyze the entire model. We conduct an ablation study on the ground truth estimation capability of our consecutive Brownian Bridge diffusion. We compare the evaluation results of decoded images with diffusion-generated latent representation $\hat{\mathbf{x}}$ and ground truth $\mathbf{x}$, which is encoded $I_{n}$. The results are shown in Table 2. It is important to note that, fixing inputs as the ground truth, our autoencoder achieves a stronger performance than the autoencoder in LDMVFI (Danier et al., 2024), indicating the effectiveness of our autoencoder. Also, fixing the autoencoder, our consecutive Brownian Bridge diffusion achieves almost identical performance with the ground truth, indicating its strong capability of ground truth estimation. However, the conditional generation model in LDMVFI (Danier et al., 2024) usually underperforms the autoencoder with ground truth inputs. Therefore, our method has a stronger ability in both the autoencoder and ground truth estimation stages. More ablation studies are provided in Appendix C.3.
@@ 5. Conclusion
In this study, we propose our consecutive Brownian Bridge diffusion Model that better estimates the ground truth latent representation due to its low cumulative variance. We justify its effectiveness with extensive experiments in a wide range of datasets, though we do acknowledge that it requires larger GPU memory (18.55G for one $1080\times 720$ image) than recent non-diffusion VFI methods such as UPR-Net (Jin et al., 2023)(3G). Our method improves when the autoencoder is improved and achieves state-of-the-art performance with a simple yet effective design of the autoencoder, demonstrating its strong potential in the VFI task as a carefully designed autoencoder could potentially boost the performance by a large margin. In addition, we propose a novel method to analyze LDM-based VFI, providing insights for future research: whether future research could be conducted on autoencoder or diffusion model. Therefore, we believe our work will provide unique research directions and insights for diffusion-based video frame interpolation.